# Imaging of I Zw 18 by JWST: I. Strategy and First Results of Dusty Stellar Populations

### v1

@@ Abstract
We present a James Webb Space Telescope (JWST) imaging survey of I Zw 18, the archetypal extremely metal-poor, star-forming, blue compact dwarf galaxy. With an oxygen abundance of only 3% , it is among the lowest-metallicity systems known in the local universe, and is, therefore, an excellent accessible analog for the galactic building blocks which existed at early epochs of ionization and star formation. These JWST data provide a comprehensive infrared (IR) view of I Zw 18 with eight filters utilizing both NIRCam (F115W, F200W, F356W, and F444W) and MIRI (F770W, F1000W, F1500W, and F1800W) photometry, which we have used to identify key stellar populations that are bright in the near- and mid-IR. These data allow for a better understanding of the origins of dust and dust-production mechanisms in metal-poor environments by characterizing the population of massive, evolved stars in the red supergiant (RSG) and asymptotic giant branch (AGB) phases. In addition, it enables the identification of the brightest dust-enshrouded young stellar objects (YSOs), which provide insight into the formation of massive stars at extremely low metallicities typical of the very early universe. This paper provides an overview of the observational strategy and data processing, and presents first science results, including identifications of dusty AGB star, RSG, and bright YSO candidates. These first results assess the scientific quality of JWST data and provide a guide for obtaining and interpreting future observations of the dusty and evolved stars inhabiting compact dwarf star-forming galaxies in the local universe.
@@ 1 Introduction
Metal-poor star-forming (SF) dwarf galaxies in the local universe represent accessible analogs to those at high redshift (e.g., Motiño Flores et al., 2021). As the enrichment history for a given system traces the build up of heavy elements from successive generations of stellar nucleosynthesis, a low-abundance galaxy mimics the astrophysical conditions common in the early universe, including the universal epoch of peak star formation (“cosmic noon”; 1.5–2; Madau & Dickinson, 2014; Van Sistine et al., 2016), at which point a majority of the universe’s star formation and chemical enrichment is expected to have taken place. At the very lowest metallicities, we may therefore approximate the SF environments of the time period shortly after the Big Bang. I Zw 18 is among the most extremely metal-poor (XMP; 0.1 ) systems known (12+(O/H) = 7.17 0.04), with a measured gas-phase oxygen abundance of only approximately 3% solar (e.g., Alloin et al., 1979; Lequeux et al., 1979; French, 1980; Kinman & Davidson, 1981; Pagel et al., 1992; Skillman & Kennicutt, 1993; Izotov et al., 1999; Lebouteiller et al., 2013). At a distance of 18.2 1.5 Mpc (Aloisi et al., 2007) and with global star-formation rate (SFR) values measured at 0.13 M yr (computed via H i absorption; Lebouteiller et al., 2013) to 0.17 M yr (estimated from radio free-free continuum emission; Hunt et al., 2005), with localized values as high as 1 M yr (via statistical analysis; Annibali et al., 2013) it is an ideal laboratory for study of both the young and evolved star demographics in an environment analogous to that found in the very early universe. Typical of such systems I Zw 18 has an enhanced total X-ray luminosity per unit of star formation rate (Lebouteiller et al., 2017). Additional properties of I Zw 18 are listed in Table 1.
@@
As the archetypal blue compact dwarf (BCD) galaxy, I Zw 18 is currently experiencing a period of intensely elevated SFR (Janowiecki & Salzer, 2014; Janowiecki et al., 2017) and has thus been the subject of focused observations since its discovery (Zwicky, 1966; Sargent & Searle, 1970; Searle & Sargent, 1972). Early descriptions of this system identified two bright regions of star formation separated by 5″.8 which constitute the main body, subsequently referred to as the northwest (NW) and southeast (SE) components. Active star formation in the brighter NW component was established via observations tracing ionized gas emission (e.g., Hunter & Thronson, 1995) and Wolf-Rayet (WR) stars (e.g., Izotov et al., 1997; Legrand et al., 1997; de Mello et al., 1998; Brown et al., 2002). This region has an H i envelope which contains inherently metal-free gas pockets (Lebouteiller et al., 2013) amongst inhomogeneities of the multi-phase interstellar medium (ISM; James et al., 2014). An additional diffuse SF region 22″ northwest of the main body (known as Component C) was identified later (Davidson et al., 1989; Dufour & Hester, 1990), with radial velocity studies of the ionized and neutral (H i) gas demonstrating its physical association and common H i envelope with I Zw 18, respectively (e.g., Dufour et al., 1996; Izotov & Thuan, 1998; van Zee et al., 1998).
@@
I Zw 18 has been intensely targeted by the Hubble Space Telescope (HST), and while initial studies found no evidence for an old, existing stellar population (e.g., Hunter & Thronson, 1995), subsequent work utilizing deeper data confirmed the presence of an early star-formation episode based on evidence of asymptotic giant branch (AGB) stars (Aloisi et al., 1999) and gas-phase abundance study (Garnett et al., 1997). Details concerning an intense, recent bout of star formation which began at least 30–50 Myr ago were published by Dufour et al. (1996), with the near-IR study of Östlin (2000) concluding that the stellar population of I Zw 18 is made up of a younger population of red supergiants (RSGs) and an older population of AGB stars, also found by Izotov & Thuan (2004). Subsequent study of these evolved stars by Östlin & Mouhcine (2005) employed HST medium-band filters to additionally identify a subset of five AGB stars via C absorption. The underlying stellar population of Component C was found to be similarly old, while exhibiting some moderate recent activity (Aloisi et al., 1999; Izotov & Thuan, 2004; Jamet et al., 2010).
@@
RA (J2000) | 09:34:02 | Izotov & Thuan (2016) |
Dec (J2000) | +55:14:28 | Izotov & Thuan (2016) |
@@
Recessional velocity | 751 2 km s | Thuan et al. (1999) |
@@
Stellar mass (M) | 2.0 10 M | Madden et al. (2014) |
Gas mass (H i) | 1.0 10 M | Lelli et al. (2012) |
Dust mass | 1.1 10 M | Herrera-Camus et al. (2012) |
@@
SFR | 0.1 – 1 M yr | Lebouteiller et al. (2013); Annibali et al. (2013) |
SFR/M (sSFR) | 10 – 10 yr | Aloisi (priv. comm.) |
@@
Establishing a census of evolved stars in I Zw 18 will help to provide for a better understanding of the origins and role of interstellar dust at the earliest epoch of our universe. While supernovae (SNe) have been found to quickly produce dust in significant quantities locally, it is unclear how similarly this mechanic behaves at the extremely low levels of metal enrichment which characterized galaxies at high redshift. In addition, the rate of both new and existing dust destruction by SNe via the passage of shocks may rival the rate of creation (Temim et al., 2015)! AGB stars are known to be a major source of dust production within nearby systems (e.g., Matsuura et al., 2009). As they evolve, thermal pulsations and convective cell activity push the products of nucleosynthetic reactions from inside the stellar core toward the stellar surface, where its ejection enriches the surrounding ISM. Because AGB stars represent the final evolutionary stage for the vast majority of stars which have left the main sequence (MS), their contributions of dust and heavy elements to the ISM due to mass loss, and the subsequent evolutionary effects on the ISM and other stars, is significant (Höfner & Olofsson, 2018).
@@
In metal-poor environments and at high redshift, the dust contribution from these AGB stars is currently not well defined. For ancient galaxies populating the early universe, prevailing expectation had been that their low- to intermediate-mass MS progenitor stars would have required a greater amount of time than had presently elapsed in order to evolve to the AGB phase, suggesting that they cannot contribute significantly to the overall dust budget. Recent work, however, has begun casting doubt on this conventional wisdom, with evidence for such evolved stars in the low-metallicity galaxy Sextans A, suggesting that these stars could contribute dust at early epochs. Additionally, observational evidence for large reservoirs of dust existing at redshifts up to 6 has been found (Bertoldi et al., 2003; Robson et al., 2004; Beelen et al., 2006; Algera et al., 2023), and perhaps even as distant as 10 (Curtis-Lake et al., 2023). Furthermore, studies of nearby low-abundance systems such as dwarf galaxies and globular clusters have shown that these environments possess dust originating from evolved stars (McDonald et al., 2010; Whitelock et al., 2018; Jones et al., 2018), however the effects of metallicity on AGB star dust production are disputed (van Loon et al., 2005; McDonald et al., 2011; Sloan et al., 2012, 2016).
@@
Observational efforts in the IR searching for metal-poor local galaxies possessing populations of evolved stars are ongoing. Prior to the advent of JWST, the state of the art for resolved extragalactic stellar populations was the Surveying the Agents of a Galaxy’s Evolution (SAGE) program for the Large (Meixner et al., 2006; Blum et al., 2006; Whitney et al., 2008; Bernard et al., 2008; Jones et al., 2015b) and Small (Gordon et al., 2011; Boyer et al., 2011; Sewiło et al., 2013) Magellanic Clouds (LMC/SMC), taken with the Spitzer Space Telescope. In addition to providing censuses of evolved stars, dust-production and mass-injection rates into the ISM were quantified (e.g., Matsuura et al., 2009; Boyer et al., 2012; Riebel et al., 2012; Srinivasan et al., 2016), with significant dust-production (10 M yr) via oxygen-rich AGB stars found in the LMC ( ). The DUST in Nearby Galaxies with Spitzer survey (DUSTiNGS; Boyer et al., 2015a, b; McQuinn et al., 2017; Boyer et al., 2017; Goldman et al., 2019a) sampled 50 additional nearby (within 1.5 Mpc) galaxies using Spitzer 3.6 and 4.5 m observations, finding that AGB stars are a major contributor of interstellar dust even at low metallicities. Recent works observing the nearby SF galaxy NGC 6822 (Lenkić et al., 2023; Nally et al., 2023) with the James Webb Space Telescope (JWST; Gardner et al., 2023) showcase its unparalleled IR sensitivity, improving upon earlier studies utilizing Spitzer and ground-based data (Jones et al., 2019; Hirschauer et al., 2020). Even with the crowded fields and large distance to I Zw 18, JWST is capable of detecting and resolving individual evolved red stars, including RSGs and both oxygen- and carbon-rich AGB stars, as well as those undergoing substantial dust loss via superwind in which the most significant amounts of dust are produced (i.e., “x-AGB” stars; Blum et al., 2006; Boyer et al., 2011, 2015a).
@@
Note. — Observations for JWST PID: 1233 executed on 27 October 2022.
JWST additionally offers incomparable sensitivity to dust-enshrouded young stellar objects (YSOs) and pre-main sequence (PMS) stars (e.g., Jones et al., 2023). The XMP abundance level, high SFR, and striking blue appearance of I Zw 18 affirms the extreme nature of its SF regions. As YSOs are born in regions of active star formation, they exhibit strong IR excess as their light is absorbed and re-emitted by their surrounding cool dusty envelopes and accretion disks. Previous photometric studies with earlier facilities were only capable of detecting the most massive such stars (8 M), which account for only roughly one in every ten thousand, and star clusters (Whitney et al., 2008; Gruendl & Chu, 2009; Carlson et al., 2012; Nayak et al., 2023). Because YSOs rapidly progress through the different stages of their evolution, observations are relatively rare. Detailed study of the population of young stars in I Zw 18 will allow for a more complete understanding of the physical mechanisms which govern star formation in environments typical of the very early universe.
@@ 2 Observing Program
We imaged I Zw 18 with JWST utilizing both the Near Infrared Camera (NIRCam; Rieke et al., 2005, 2023) and the Mid-Infrared Instrument (MIRI; Rieke et al., 2015; Wright et al., 2023) imager (Bouchet et al., 2015, Dicken et al. submitted). In total, 6.49 hours of integration time were split between NIRCam (1.91 hours) and MIRI (4.58 hours), across eight filters (four filters each) employed for the prime observations. The observations with NIRCam were taken on 27 October 2022, utilizing the F115W and F200W short-wavelength (SW) filters, alongside the F356W and F444W long-wavelength (LW) filters, with I Zw 18 centered in the upper-left quadrant of Module B (B3), and Module A observing empty sky to provide suitable data for background contamination subtraction. Employing both the A and B modules with a single pointing, these observations covered a total field of view (FoV) of 9.7 arcmin, with two 2.2′2.2′ fields separated by a 44″ gap.
@@
The spatial coverage for the two instruments is illustrated in Figure 1, and principal characteristics of this program are summarized in Table 2. At a distance of 18.2 Mpc (Aloisi et al., 2007), the angular resolution is 0.04 – 0.14″ (3.5 – 12.4 pc) in the NIRCam filters and 0.3 – 0.6″ (26.5 – 53.1 pc) in the MIRI filters, respectively. This provides an improvement over existing IR imaging with Spitzer (spatial resolution 2″) by greater than an order of magnitude, and is of similar quality to optical-band imaging with HST.

### v2

# Imaging of I Zw 18 by JWST. I. Detecting Dusty Stellar Populations

@@ Abstract
We present a JWST imaging survey of I Zw 18, the archetypal extremely metal-poor, star-forming (SF), blue compact dwarf galaxy. With an oxygen abundance of only 3% , it is among the lowest-metallicity systems known in the local Universe, and is, therefore, an excellent accessible analog for the galactic building blocks which existed at early epochs of ionization and star formation. These JWST data provide a comprehensive infrared (IR) view of I Zw 18 with eight filters utilizing both Near Infrared Camera (F115W, F200W, F356W, and F444W) and Mid-Infrared Instrument (F770W, F1000W, F1500W, and F1800W) photometry, which we have used to identify key stellar populations that are bright in the near- and mid-IR. These data allow for a better understanding of the origins of dust and dust-production mechanisms in metal-poor environments by characterizing the population of massive, evolved stars in the red supergiant (RSG) and asymptotic giant branch (AGB) phases. In addition, it enables the identification of the brightest dust-enshrouded young stellar objects (YSOs), which provide insight into the formation of massive stars at extremely low metallicities typical of the very early Universe. This paper provides an overview of the observational strategy and data processing, and presents first science results, including identifications of dusty AGB, RSG, and bright YSO candidates. These first results assess the scientific quality of JWST data and provide a guide for obtaining and interpreting future observations of the dusty and evolved stars inhabiting compact dwarf SF galaxies in the local Universe.
@@ 1 Introduction
Metal-poor star-forming (SF) dwarf galaxies in the local Universe represent accessible analogs to those at high redshift (e.g., Motiño Flores et al., 2021). As the enrichment history for a given system traces the buildup of heavy elements from successive generations of stellar nucleosynthesis, a low-abundance galaxy mimics the astrophysical conditions common in the early Universe, including the universal epoch of peak star formation (“cosmic noon,” 1.5–2; Madau & Dickinson, 2014; Van Sistine et al., 2016), at which point a majority of the Universe’s star formation and chemical enrichment is expected to have taken place. At the very lowest metallicities, we may therefore approximate the SF environments of the time period shortly after the Big Bang. I Zw 18 is among the most extremely metal-poor (XMP; 0.1 ) systems known (12 + (O/H)=7.17 0.04), with a measured gas-phase oxygen abundance of only approximately 3% solar (e.g., Alloin et al., 1979; Lequeux et al., 1979; French, 1980; Kinman & Davidson, 1981; Pagel et al., 1992; Skillman & Kennicutt, 1993; Izotov et al., 1999; Lebouteiller et al., 2013). At a distance of 18.2 1.5 Mpc (Aloisi et al., 2007) and with global star-formation rate (SFR) values measured at 0.13 yr-1 (computed via H I absorption; Lebouteiller et al., 2013) to 0.17 yr-1 (estimated from radio free-free continuum emission; Hunt et al., 2005), with localized values as high as 1 yr-1 (via statistical analysis; Annibali et al., 2013) it is an ideal laboratory for study of both the young and evolved star demographics in an environment analogous to that found in the very early Universe. Typical of such systems, I Zw 18 has an enhanced total X-ray luminosity per unit of SFR (Lebouteiller et al., 2017). Additional properties of I Zw 18 are listed in Table 1.
@@
As the archetypal blue compact dwarf (BCD) galaxy, I Zw 18 is currently experiencing a period of intensely elevated SFR (Janowiecki & Salzer, 2014; Janowiecki et al., 2017) and has thus been the subject of focused observations since its discovery (Zwicky, 1966; Sargent & Searle, 1970; Searle & Sargent, 1972). Early descriptions of this system identified two bright regions of star formation separated by 5″.8 which constitute the main body, subsequently referred to as the northwest (NW) and southeast (SE) components. Active star formation in the brighter NW component was established via observations tracing ionized gas emission (e.g., Hunter & Thronson, 1995) and Wolf–Rayet (WR) stars (e.g., Izotov et al., 1997; Legrand et al., 1997; de Mello et al., 1998; Brown et al., 2002). This region has an H I envelope which contains inherently metal-free gas pockets (Lebouteiller et al., 2013) amongst inhomogeneities of the multiphase interstellar medium (ISM; James et al., 2014). An additional diffuse SF region 22″ NW of the main body (known as Component C) was identified later (Davidson et al., 1989; Dufour & Hester, 1990), with radial velocity studies of the ionized and neutral (H I) gas demonstrating its physical association and common H I envelope with I Zw 18, respectively (e.g., Dufour et al., 1996; Izotov & Thuan, 1998; van Zee et al., 1998).
@@
I Zw 18 has been intensely targeted by the Hubble Space Telescope (HST), and while initial studies found no evidence for an old, existing stellar population (e.g., Hunter & Thronson, 1995), subsequent work utilizing deeper data confirmed the presence of an early star-formation episode based on evidence of asymptotic giant branch (AGB) stars (Aloisi et al., 1999) and gas-phase abundance study (Garnett et al., 1997). Details concerning an intense, recent bout of star formation which began at least 30–50 Myr ago were published by Dufour et al. (1996), with the near-infrared (near-IR) study of Östlin (2000) concluding that the stellar population of I Zw 18 is made up of a younger population of red supergiants (RSGs) and an older population of AGB stars, also found by Izotov & Thuan (2004). Subsequent study of these evolved stars by Östlin & Mouhcine (2005) employed HST medium-band filters to additionally identify a subset of five AGB stars via C2 absorption. The underlying stellar population of Component C was found to be similarly old, while exhibiting some moderate recent activity (Aloisi et al., 1999; Izotov & Thuan, 2004; Jamet et al., 2010).
@@
| R.A. (J2000) | 09:34:02 | Izotov & Thuan (2016) |
| Decl. (J2000) | +55:14:28 | Izotov & Thuan (2016) |
@@
| Recessional velocity | 751 2 km s-1 | Thuan et al. (1999) |
@@
| Stellar mass () | 2.0 107 M⊙ | Madden et al. (2014) |
| Gas mass (H I) | 1.0 108 M⊙ | Lelli et al. (2012) |
| Dust mass | 1.1 104 M⊙ | Herrera-Camus et al. (2012) |
@@
| SFR | 0.1– 1 M⊙ yr-1 | Lebouteiller et al. (2013); Annibali et al. (2013) |
| SFR/ (sSFR) | 10-7–10-8 yr-1 | A. Aloisi (2024, private communication) |
@@
Establishing a census of evolved stars in I Zw 18 will help to provide for a better understanding of the origins and role of interstellar dust at the earliest epoch of our Universe. While supernovae (SNe) have been found to quickly produce dust in significant quantities locally, it is unclear how similarly this mechanic behaves at the extremely low levels of metal enrichment which characterized galaxies at high redshift. In addition, the rate of both new and existing dust destruction by SNe via the passage of shocks may rival the rate of creation (Temim et al., 2015)! AGB stars are known to be a major source of dust production within nearby systems (e.g., Matsuura et al., 2009). As they evolve, thermal pulsations and convective cell activity push the products of nucleosynthetic reactions from the internal regions previously exposed to helium burning reactions towards toward the stellar surface (e.g., Iben, 1974, 1975; Lattanzio, 1987, 1993; Boothroyd & Sackmann, 1988), where its ejection enriches the surrounding ISM. Because AGB stars represent the final evolutionary stage for the vast majority of stars which have left the main sequence (MS), their contributions of dust and heavy elements to the ISM due to mass loss, and the subsequent evolutionary effects on the ISM and other stars, is significant (Höfner & Olofsson, 2018).
@@
In metal-poor environments and at high redshift, the dust contribution from these AGB stars is currently not well defined. For ancient galaxies populating the early Universe, the prevailing expectation had been that their low- to intermediate-mass MS progenitor stars would have required a greater amount of time than had presently elapsed in order to evolve to the AGB phase, suggesting that they cannot contribute significantly to the overall dust budget. Recent work, however, has begun casting doubt on this conventional wisdom, with evidence for such evolved stars in the low-metallicity galaxy Sextans A suggesting that these stars could contribute dust at early epochs. Additionally, observational evidence for large reservoirs of dust existing at redshifts up to 6 has been found (Bertoldi et al., 2003; Robson et al., 2004; Beelen et al., 2006; Algera et al., 2023), and perhaps even as distant as 10 (Curtis-Lake et al., 2023). Furthermore, studies of nearby low-abundance systems such as dwarf galaxies and globular clusters have shown that these environments possess dust originating from evolved stars (McDonald et al., 2010; Whitelock et al., 2018; Jones et al., 2018); however, the effects of metallicity on AGB star dust production are disputed (van Loon et al., 2005; McDonald et al., 2011; Sloan et al., 2012, 2016; Dell’Agli et al., 2019b).
@@
Observational efforts in the IR searching for metal-poor local galaxies possessing populations of evolved stars are ongoing. Prior to the advent of JWST, the state of the art for resolved extragalactic stellar populations was the Surveying the Agents of a Galaxy’s Evolution (SAGE) program for the Large (Meixner et al., 2006; Blum et al., 2006; Whitney et al., 2008; Bernard et al., 2008; Jones et al., 2015b) and Small (Gordon et al., 2011; Boyer et al., 2011; Sewiło et al., 2013) Magellanic Clouds (LMC/SMC), taken with the Spitzer Space Telescope (Spitzer). In addition to providing censuses of evolved stars, dust-production and mass-injection rates into the ISM were quantified (e.g., Matsuura et al., 2009; Boyer et al., 2012; Riebel et al., 2012; Srinivasan et al., 2016), with significant dust production (10-4 yr-1) via oxygen-rich AGB stars found in the LMC ( ). The DUST in Nearby Galaxies with Spitzer survey (DUSTiNGS; Boyer et al., 2015a, b; McQuinn et al., 2017; Boyer et al., 2017; Goldman et al., 2019a) sampled 50 additional nearby (within 1.5 Mpc) galaxies using Spitzer 3.6 and 4.5 m observations, finding that AGB stars are a major contributor of interstellar dust even at low metallicities. Recent works observing the nearby SF galaxy NGC 6822 (Lenkić et al., 2024; Nally et al., 2024) with the JWST (Gardner et al., 2023) showcase its unparalleled IR sensitivity, improving upon earlier studies utilizing Spitzer and ground-based data (Jones et al., 2019; Hirschauer et al., 2020). Even with the crowded fields and large distance to I Zw 18, JWST is capable of detecting and resolving individual evolved red stars, including RSGs and both oxygen- and carbon-rich AGB stars, as well as those undergoing substantial dust loss via superwind in which the most significant amounts of dust are produced (i.e., “x-AGB” stars; Blum et al., 2006; Boyer et al., 2011, 2015a).
@@
Note. — Observations for JWST PID: 1233 executed on 2022 October 27.
JWST additionally offers incomparable sensitivity to dust-enshrouded young stellar objects (YSOs) and pre-main-sequence (PMS) stars (e.g., Jones et al., 2023). The XMP abundance level, high SFR, and striking blue appearance of I Zw 18 affirms the extreme nature of its SF regions. As YSOs are born in regions of active star formation, they exhibit strong IR excess as their light is absorbed and reemitted by their surrounding cool dusty envelopes and accretion disks. Previous photometric studies with earlier facilities were only capable of detecting the most massive such stars (8), which account for only roughly one in every 10,000, and star clusters (Whitney et al., 2008; Gruendl & Chu, 2009; Carlson et al., 2012; Nayak et al., 2023). Because YSOs rapidly progress through the different stages of their evolution, observations are relatively rare. Detailed study of the population of young stars in I Zw 18 will allow for a more complete understanding of the physical mechanisms which govern star formation in environments typical of the very early Universe.
@@ 2 Observing Program
We imaged I Zw 18 with JWST utilizing both the Near Infrared Camera (NIRCam; Rieke et al., 2005, 2023) and the Mid-Infrared Instrument (MIRI; Rieke et al., 2015; Wright et al., 2023) imager (Bouchet et al., 2015; Dicken et al., 2024). In total, 6.49 hr of integration time were split between NIRCam (1.91 hr) and MIRI (4.58 hr), across eight filters (four filters each) employed for the prime observations. The observations with NIRCam were taken on 2022 October 27, utilizing the F115W and F200W short-wavelength (SW) filters, alongside the F356W and F444W long-wavelength (LW) filters, with I Zw 18 centered in the upper-left quadrant of Module B (B3), and Module A observing empty sky to provide suitable data for background contamination subtraction. Employing both the A and B modules with a single pointing, these observations covered a total field of view (FOV) of 9.7 arcmin2, with two 2′.2 2′.2 fields separated by a 44″ gap.
@@
The spatial coverage for the two instruments is illustrated in Figure 1, and the principal characteristics of this program are summarized in Table 2. At a distance of 18.2 Mpc (Aloisi et al., 2007), the angular resolution is 0″.04 – 0″.14 (3.5 – 12.4 pc) in the NIRCam filters and 0″.3 – 0″.6 (26.5 – 53.1 pc) in the MIRI filters, respectively. This provides an improvement over existing IR imaging with Spitzer (spatial resolution 2″) by greater than an order of magnitude, and is of similar quality to optical-band imaging with HST.